Prove each case f(n) = O(g(n)) by providing enough justification.

1) f(n) = n2.5

2) f(n) = n10
Answer #1
Big-Oh definition:
-------------------
f(n) = O(g(n)) means there are positive constants c and k, such that 0 ≤ f(n) ≤ cg(n) for all n ≥ k.

1) f(n) = n2.5
0 <= f(n) <= 2n^2.5
Where c=2, n>0 and g(n)=n^2.5
So, from the definition of Big-Oh we can say that f(n)=

2) f(n) = n10
0 <= f(n) <= 2n^10
Where c=2, n>0 and g(n)=n^10
So, from the definition of Big-Oh we can say that f(n)=